OpenStudy (anonymous):

Calculate the result of the following formula: P(14,7) Don't want the answer, just don't understand how to solve.

ganeshie8 (ganeshie8):

\[\rm P(n, r) = \dfrac{n!}{(n-r)!}\]

ganeshie8 (ganeshie8):

\[\rm P(14, 7) = \dfrac{14!}{(14-7)!}\]

ganeshie8 (ganeshie8):

do you know how to work a factorial ?

OpenStudy (anonymous):

We eliminate the the overlapping numbers to reduce the work I think...

OpenStudy (anonymous):

14*13*12...*3*2*1

ganeshie8 (ganeshie8):

yes!

OpenStudy (anonymous):

The number seems awfully large which is throwing me for a loop.

ganeshie8 (ganeshie8):

\[\rm P(14, 7) = \dfrac{14!}{(14-7)!}= \dfrac{14!}{7!}\\\]

ganeshie8 (ganeshie8):

\[ = \dfrac{14\times 13\times 12\times 11\times 10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2\times 1}{7\times 6\times 5\times 4\times 3\times 2\times 1}\]

ganeshie8 (ganeshie8):

\[\require{cancel} = \dfrac{14\times 13\times 12\times 11\times 10\times 9\times 8\times \cancel{7\times 6\times 5\times 4\times 3\times 2\times 1}}{\cancel{7\times 6\times 5\times 4\times 3\times 2\times 1}^1}\]

OpenStudy (anonymous):

Ohhhhhh!!!!!

OpenStudy (anonymous):

I KNEW there was some sort of elimination process.

ganeshie8 (ganeshie8):

\[=14\times 13\times 12\times 11\times 10\times 9\times 8 \]

ganeshie8 (ganeshie8):

:) you still need use your calculator ^

OpenStudy (anonymous):

Yea, it's still a giant number.

OpenStudy (anonymous):

So I have this oversized number. What's to be done with it now?

ganeshie8 (ganeshie8):

you're done, thats the value of P(14, 7)

OpenStudy (anonymous):

p(14,7)=121080960

OpenStudy (anonymous):

That just seems...excessive.

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=P%2814%2C7%29

OpenStudy (anonymous):

How'd I goof THAT up...

ganeshie8 (ganeshie8):

looks you have multiplied by 7

ganeshie8 (ganeshie8):

you should stop at 8

OpenStudy (anonymous):

Yep. Spot on. Do you, by chance, know the name of that formula?

ganeshie8 (ganeshie8):

P(14,7) = 14x13x12x11x10x9x8

ganeshie8 (ganeshie8):

permutation formula

OpenStudy (anonymous):

You're top notch. Thanks so much for help! I have a much better grasp on these sort of equations and that formula.

ganeshie8 (ganeshie8):

you're welcome :)